Approximately how many times will a 50,000 investment, earning 5.54% interest, double from age 25 to age 64? what will the investment be worth?

Please show work so I can understand.
Answer #1

There are two types calculation based on the types of interest.

a) Simple Interest & b) Compound interest

A) Simple Interest of 5.54%

Interest = Principal X Interest rate X Number of years

Interest = 50000 x 5.54% x 39 years = 108030

Number of times = 108030 / 50000 = 2.16 times

The interest received is 2.16 times the principal amount

B) Compound Interest @ 5.54% compounding annually

Amount = Principal X (1 + Interest rate) ^ (number of years) Amount = 50000 X (1 + 0.0554) ^ (39) = 50000 x 8.1897

Amount = 409484

Interest= 409484-50000 = 359484.

No. Of times = 359484/50000 = 7.1897 times

Interest received is 7.1897 times of the principal